OpenStudy (anonymous):

Please help? write the equations for tangent with the period=1/3pi, phase shift=-1/4pi, vertical shift= -5 y=tan(3theta+3/4pi)-5 y=tan(1/3theta+3/4pi)+5 y=tan(1/3theta-3/4pi)+5 y=tan(3theta-3/4pi)-5

OpenStudy (shamil98):

For F(t) = A(Bt – C) + D, where f(t) is one of the basic trig functions, we have: A: amplitude is A B: period is (2π)/|B| C: phase shift is C/B D: vertical shift is D Identify what you have.

OpenStudy (shamil98):

y = tan ( 1/3 pi + 1/4 pi ) - 5

OpenStudy (shamil98):

Oh i made an error

OpenStudy (shamil98):

-1/4 x 3/1 = phase shift = -3/4

OpenStudy (anonymous):

Wow, thanks!!

OpenStudy (shamil98):

y = Atan(Bt – C) + D y = 1 tan ( 1/3 pi - (-3/4) + (-5)

OpenStudy (shamil98):

y = tan (1/3pi + 3/4) - 5

OpenStudy (shamil98):

My trig is somewhat rusty :S

OpenStudy (anonymous):

Thanks for all the help! I think I understand it now ^^

OpenStudy (shamil98):

and I forgot the pi symbol afterwards again lol y = tan ( 1/3 pi + 3/4 pi ) - 5 Yeah, no problem.

OpenStudy (shamil98):

If you need more help in understanding this, I know it may be long but it helps http://www.purplemath.com/modules/grphtrig.htm

OpenStudy (anonymous):

Okay, thanks ^^ My text didnt go into much detail so I this'll make more sense.

OpenStudy (shamil98):

This one is a bit more detailed and easier. http://www.regentsprep.org/regents/math/algtrig/ATT7/indexATT7.htm